### v1

@@ Abstract
Interdisciplinary collaboration is crucial for addressing complex scientific challenges. Recent advancements in large language models (LLMs) have shown significant potential in benefiting researchers across various fields. To explore the application of LLMs in scientific disciplines and their implications for interdisciplinary collaboration, we collect and analyze 50,391 papers from OpenAlex, an open-source platform for scholarly metadata. We first employ Shannon entropy to assess the diversity of collaboration in terms of authors’ institutions and departments. Our results reveal that most fields have exhibited varying degrees of increased entropy following the release of ChatGPT, with Computer Science displaying a consistent increase. Other fields such as Social Science, Decision Science, Psychology, Engineering, Health Professions, and Business, Management & Accounting have shown minor to significant increases in entropy in 2024 compared to 2023. Statistical testing further indicates that the entropy in Computer Science, Decision Science, and Engineering is significantly lower than that in health-related fields like Medicine and Biochemistry, Genetics & Molecular Biology. In addition, our network analysis based on authors’ affiliation information highlights the prominence of Computer Science, Medicine, and other Computer Science-related departments in LLM research. Regarding authors’ institutions, our analysis reveals that entities such as Stanford University, Harvard University, University College London, and Google are key players, either dominating centrality measures or playing crucial roles in connecting research networks. Overall, this study provides valuable insights into the current landscape and evolving dynamics of collaboration networks in LLM research. Our findings also suggest potential areas for fostering more diverse collaborations and highlight the need for continued research on the impact of LLMs on scientific research practices and outcomes.
@@
Large language model, interdisciplinary collaboration, diversity analysis, Shannon entropy, network analysis
@@ 1 Introduction
Interdisciplinary collaboration is increasingly recognized as critical for addressing complex scientific challenges, which often involves investigators from multiple fields of expertise [1, 2, 3]. Recent advances in generative AI models and applications, particularly the development of large language models (LLM) such as OpenAI’s ChatGPT, Google’s PaLM, and Meta’s Llama, are reshaping research activities across disciplines. Researchers have demonstrated LLMs’ utility in research tasks such as literature search [4, 5], content analysis [6], and findings generation from provided data [7].
@@
The impact of LLMs extends far beyond their computer science origins or general application in research activities. In health-related fields, LLMs are being employed to interpret protein structures [8], process electronic health records [9], and even aid in drug discovery [10]. Engineering benefits from these models through advancements in autonomous driving [11] and remote sensing [12] technologies. Social scientists are leveraging LLMs for large-scale text analysis, including attitude simulation [13] and content moderation on social media platforms [14]. Their impact also extends to finance, where they’re being used to streamline document review and perform financial analysis [15]. This wide-ranging applicability across disciplines drives us to believe that LLMs are transforming multiple aspects of the research processes, potentially impacting how researchers collaborate.
@@
Traditionally, interdisciplinary collaboration has been essential to combine expertise from diverse disciplines to find novel solutions to scientific problems. However, the advent of LLMs is reshaping this paradigm. Researchers have leveraged tools like ChatGPT to gain cross-disciplinary insights and interpret their work from various perspectives, potentially reducing the need for direct collaboration with experts from other fields [16]. LLMs have also been used to automate certain aspects of scholarly writing process, such as synthesizing existing literature from multiple disciplines on a specific topic [17], and reframing the main findings from different disciplines into domain-specific language [18].
@@
While LLMs may have lowered the barriers to conducting interdisciplinary research, especially for those unfamiliar with language processing techniques, they offer significant potential to enhance research practices. By automating routine tasks [5, 19], such as crunching numbers and formatting manuscripts according to journal guidelines, LLMs can free up time for researchers to focus on developing innovative solutions and gaining deeper insights [20, 17]. Well-known LLMs such as ChatGPT have also illustrated capabilities to bridge language gaps between disciplines such as environmental science and ecology [17], statistics and biology [18]. However, it’s crucial to recognize that LLMs’ effectiveness is primarily illustrated in language-related tasks. They are not substitutes for the critical examination of findings or the nuanced and domain-specific understanding essential for truly bridging disciplinary language [21, 22].
@@
Given that the use of LLMs is reshaping the way scholars work and collaborate, the questions then arise: how have LLMs transformed interdisciplinary collaboration? In what ways have they changed how different disciplines interact and collaborate? Therefore, our study explores the application of LLMs in scientific disciplines and their implications for interdisciplinary collaboration by addressing two key research questions, as listed below. The first question aims to examine the diversity of co-authors in terms of their institutional and departmental affiliations, using the Shannon entropy measure. The second question aims to understand the structural patterns of co-authorship collaborations, using network analysis to identify key researchers, institutions, and disciplines that are active producers of LLMs research, as well as those facilitating collaborations across disciplinary boundaries.
@@
  • RQ1. How diverse are the coauthors of papers utilizing LLMs in terms of their research institutions and departments across different fields?
@@
  • RQ2. What are the structural patterns of co-authorship networks in research utilizing LLMs, and what roles do key entities (leading researchers, institutions, and departments) play in facilitating and enhancing collaboration?
@@
To answer these questions, we collect 50,391 papers from OpenAlex related to LLMs from 27 scientific fields published between the release of the BERT model in October 2018 and June 2024. We use Shannon entropy to measure the authorship diversity and social network techniques to reveal the collaboration network structures. Our findings indicate that since the advent of LLMs —particularly with the release of ChatGPT in November 2022—collaboration diversity has increased across many fields, notably Computer Science, Social Science, and Psychology. Medicine is the only discipline where diversity shows a significant decrease. Our network analysis further reveals that Computer Science and Medicine consistently play influential roles in connecting different research communities regarding LLM applications. Overall, our findings demonstrate that LLM research has grown exponentially since 2022 and holds the potential to enhance interdisciplinary collaboration. It also encourages the involvement of prominent academic institutions in leading LLM-based research and applications across different domains.
@@ 2 Data and Methods
Aligned with the two research questions, our data collection and methods involve two units of analysis: papers and authors. The first research question examines papers to analyze the diversity of coauthors across research fields and institutions. The second research question centers on authors to explore the structural patterns of co-authorship networks and identify key researchers and their roles in facilitating collaboration. Combining these two units of analysis allows us to capture (1) which disciplines and institutions are most impacted by the advent of LLMs in terms of collaboration diversity, as well as (2) key researchers in LLM research and their respective departmental and institutional affiliations. These analyses help us understand how LLMs are impacting scientific collaboration by revealing patterns of interdisciplinary and cross-institutional partnerships, as well as identifying the influential actors driving LLM research across disciplinary boundaries.
@@ 2.1 Data preparation
The workflow for data cleaning is presented in subsection A.1. Within OpenAlex, we collect relevant papers on the most popular LLMs and their respective models, as detailed in Table 2 in subsection A.2. We use two general terms, “large language model” and “LLM,” along with popular open-source models (e.g., BERT, Flan-T5, LLaMA) and closed-source models (e.g., ChatGPT, Claude) based on the MMLU benchmark [24]. We observe that some models, like Yi or Phi, do not yield relevant papers, potentially introducing significant noise during paper screening. Additionally, models like grok-1 or Galactic did not return any search results. Our data collection, conducted as of June 22, 2024, results in a total of 97,242 papers.
To ensure the collected papers are relevant to the topic of LLMs, we restrict our search to titles and abstracts. However, some papers containing these keywords might still be irrelevant. Therefore, we implement several steps to filter out irrelevant papers. First, we consider only articles and preprints, excluding types such as editorials and opinions. Second, we remove potential duplicates, including those with duplicated titles and papers initially published as preprints and later as journal articles with similar titles. We use Jaccard similarity to identify and remove these duplicates (see subsection A.3). Third, we employ GPT-4 models to evaluate the relevance of a paper to LLM topics based on its title, abstract, and keywords, and filter out those papers that are not relevant to LLM topics (see subsection A.4). Applying these filtering criteria reduces the original collection to 50,391 papers for the subsequent analysis.
@@ 2.2 Measure of collaboration diversity
where $x_{i}$ denotes the $i$th author of a paper, $A(x_{i})$ denotes the set of authors’ affiliation information given a paper, $D(x_{i})$ represents their department information, $I(x_{i})$ represents their institution information, and $C(x_{i})$ represents their country information (see Appendix D for the analysis of countries). It is important to note that these three types of information can vary significantly within a single paper. For instance, all collaborating authors could belong to the same institution and country but be affiliated with different departments. Our subsequent analysis particularly focuses on the collaboration between institutions and departments; therefore, we consider the first two sets of variables in Equation 1.
@@
2.3 Measures of network structure
We construct co-authorship networks based on the bipartite network projection, which involves converting a paper-author network to a co-authorship network whereby two authors are connected if they have co-authored at least one paper together. Each connection is weighted based on the total number of papers that each pair of researchers co-authored together. This method, as described by [25, 2], allows us to identify key researchers that have notable collaborative influence in the field, as well as any differences in collaboration patterns depending on the authors’ disciplines, institutions, and countries. The formula for our weighted projection approach is shown below:
@@ 3 Results
In subsection 3.1, we analyze the diversity of collaboration based on the institutions and departments affiliated with authors across various fields. In subsection 3.2, we examine the network to understand how different institutions or departments collaborate with each other. It should be noted that the field categorization of a paper is provided by OpenAlex (see Appendix C for more details). OpenAlex has established a BERT model that generates a score distribution for identified topics using a paper’s title, abstract, and citations [23]. Their model provides up to three topics with the highest scores, which are then mapped to fields according to the ASJC structure [23].
@@
In addition, we sort the collected 50,391 papers based on the paper count in order, as presented in Figure 1(a). For the subsequent analysis, we focus on the top 12 fields with the most publications in the topics of LLMs: (1) Computer Science, (2) Medicine, (3) Social Science, (4) Decision Science, (5) Biochemistry, Genetics & Molecular Biology, (6) Psychology, (7) Engineering, (8) Health Professions, (9) Business, Management & Accounting, (10) Neuroscience, (11) Arts & Humanities, and (12) Materials Science. We then filter for papers with complete authors’ affiliation information, resulting in 25,933 papers with complete institution information and 16,645 papers with complete department information. Overall, the entropy shows an increase based on authors’ affiliated institution ($coefficient=0.73$) and department ($coefficient=0.81$) information (see Figure 1(b)).
@@
Figure 1: (a) The distribution of papers in the collection by field. (b) The temporal pattern of entropy and paper count based on authors’ affiliated institution (top) and department (bottom) information, respectively. In both subplots, the x-axis denotes the date by year. The primary y-axis denotes the Shannon Entropy calculated using Equation 2, while the secondary y-axis indicates the count of papers. Each point represents the averaged entropy based on papers published in one quarter. The dotted gray line implies the date when ChatGPT was released by OpenAI.
3.1 Analysis of collaboration diversity
Figure 2(a) and Figure 2(b) show the collaboration diversity measured by Shannon Entropy for authors’ institutions and departments, respectively. It is interesting to notice that nearly all fields exhibit a sharp increasing trend after the release of ChatGPT. Before that date, only Computer Science shows a comparatively larger number of LLM-related studies. This is possible because BERT models, such as DeBERTa and RoBERTa, were widely studied before the release of ChatGPT. It is also noteworthy that other fields, such as Biochemistry, Genetics & Molecular Biology, Psychology, Neuroscience, Social Science, Engineering, and Arts & Humanities, demonstrated significant applications of LLMs even before the release of ChatGPT. After the release of ChatGPT, one of the most interesting observations regarding the trend is that Medicine shows a sharp and significant increase, with the peak close to that of Computer Science. Similarly, Business, Management & Accounting and health Professions have also witnessed a sharp and increasing trend after the release of ChatGPT.
Regarding the entropy analysis based on institutions and departments, the Shannon Entropy trends in Figure 2(a) and Figure 2(b) closely align for each respective field. Compared to 2023, most fields show a minor increase in the averaged entropy in 2024. Exceptions include Biochemistry, Genetics & Molecular Biology, Engineering, Neuroscience, Business, Management, & Accounting, and Materials Science. Moreover, the variance of entropy is much wider before ChatGPT’s release, possibly due to fewer papers on the topic of LLMs and more diverse collaborations. One possible explanation for this pattern is that researchers from other fields might have sought partnerships with Computer Science experts for such research topics.
In addition, the observed changes in entropy suggest a potential shift in research collaboration dynamics following the introduction of ChatGPT across several fields. First, papers in the field of Computer Science display a stable increasing trend. A plausible explanation for this is that many Computer Science researchers are seeking collaborations with domain experts, focusing on areas such as AI for science applications or AI for social good. However, Medicine exhibits a notable decrease in entropy. This is the only field showing a statistically significant decrease in entropy after ChatGPT’s release, while other fields display either minor increases or negligible changes. These findings highlight the varying impacts of ChatGPT’s introduction on interdisciplinary collaborations across different academic fields.
Then, we conduct a Wilcoxon rank-sum test to compare the entropy before and after the release of ChatGPT, with detailed results presented in Appendix B (see Figure 5). Our analysis reveals statistically significant changes in entropy across several fields. Computer Science, Biochemistry, Genetics & Molecular Biology, and Psychology show a statistically significant increase in entropy, while Medicine displays a statistically significant decrease. The other fields do not show any significant increase or decrease in the entropy. In particular, we observe that the mean of the entropy distribution for most fields remains at 0, with the first quartile consistently at 0. This suggests that many researchers primarily collaborate with colleagues from their own institutions or departments on LLM-related topics.
We conduct additional Wilcoxon rank-sum tests to compare entropy across different fields, with results presented as heatmaps in Figure 6. Several consistent patterns emerge when examining both institution and department results before and after ChatGPT’s release. Certain fields like Computer Science, Medicine, and Neuroscience exhibit more significant differences compared to other fields, as evidenced by darker blue cells in the heatmaps. Following ChatGPT’s release, there is also an increase in significant differences across fields, illustrated by a higher prevalence of darker cells in the right heatmap in Figure 6, compared to their counterparts in the left heatmap of the same figures. Another interesting observation is that Computer Science consistently demonstrates significantly lower entropy than Medicine, Biochemistry, Genetics & Molecular Biology, Health Professions, and Neuroscience.

### v2

@@ Abstract
Interdisciplinary collaboration is crucial for addressing complex scientific challenges. Recent advancements in large language models (LLMs) have shown significant potential in benefiting researchers across various fields. To explore the application of LLMs in scientific disciplines and their potential for interdisciplinary collaboration, we collect and analyze data from OpenAlex, an open-source academic database. Our dataset comprises 59,293 LLM-related papers, along with two control groups—a random sample of 70,945 machine learning (ML) papers and 73,110 papers from non-LLM/ML fields. We first employ Shannon Entropy to assess the diversity of collaboration in terms of authors’ institutions and departments. Our results reveal that many fields have exhibited a more significant increasing trend in entropy following the release of ChatGPT as compared to the control paper groups. In particular, Computer Science and Social Science display a consistent increase in both institution and department entropy. Other fields such as Decision Science, Psychology, and Health Professions have shown minor to significant increases in 2024 compared to 2023. Our difference-in-difference analysis also indicates that the release of ChatGPT leads to a statistically significant increase in collaboration in several fields, such as Computer Science and Social Science. In addition, we analyze the author networks and find that Computer Science, Medicine, and other Computer Science-related departments are the most prominent in LLM research. Regarding authors’ institutions, our analysis reveals that entities such as Stanford University, Harvard University, and University College London are key players, either dominating centrality measures or playing crucial roles in connecting research networks. Overall, this study provides valuable information on the current landscape and evolving dynamics of collaboration networks in LLM research. It also suggests potential areas for fostering more diverse collaborations and highlights the need for continued research on the impact of LLMs on scientific practices and outcomes.
@@
Large language model, interdisciplinary collaboration, scientific community, Shannon Entropy, network analysis
@@ 1 Introduction
Interdisciplinary collaboration is crucial for tackling complex scientific challenges, involving investigators from various fields of expertise [1, 2, 3]. Recent advances in generative AI models and applications, particularly the development of large language models (LLM) such as OpenAI’s ChatGPT, Google’s PaLM, and Meta’s Llama, impact research activities across disciplines. These models find applications in research tasks such as literature search [4, 5], content analysis [6], and findings generation from provided data [7]. The integration of LLMs into research workflows has enabled researchers to process and analyze vast amounts of scientific data more efficiently, identify patterns and connections across disciplines, and generate preliminary insights that can guide further investigation.
@@
While researchers initially developed and applied LLMs within computer science, their use has expanded into other disciplines, leading to an increasing number of studies in areas such as medicine, social sciences, and business. In health-related fields, LLMs are being employed to interpret protein structures [8], process electronic health records [9], and even aid in drug discovery [10]. Engineering benefits from these models through advancements in autonomous driving [11] and remote sensing [12] technologies. Social scientists leverage LLMs for large-scale text analysis, including attitude simulation [13] and content moderation on social media platforms [14]. Their impact also extends to finance, where professionals use LLMs to streamline document review and perform financial analysis [15]. The wide-ranging applicability of LLMs across disciplines suggests that they may influence many aspects of the research process, including how researchers collaborate.
@@
Traditionally, interdisciplinary collaboration has played a crucial role in combining expertise from various fields to discover innovative solutions to scientific challenges. Researchers increasingly use tools like ChatGPT to gain cross-disciplinary insights and interpret their work through different domains and perspectives [16]. Researchers also use LLMs to automate certain aspects of the scholarly writing process. They synthesize existing literature from multiple disciplines on a specific topic [17] and reframe the main findings into domain-specific language [18].
@@
By automating routine research tasks [5, 19] associated with manuscript formatting or data entry, LLMs can help free up time for researchers to build their collaborative networks and deepen their engagement in cross-disciplinary literature [20, 17]. Well-known LLMs such as ChatGPT have also illustrated capabilities to bridge language gaps between disciplines such as environmental science and ecology [17], statistics and biology [18]. Prior analysis of ChatGPT’s use in environmental sciences shows that LLMs may improve shared language between researchers [20], which in turn facilitates shared understanding across different domains. However, it’s important to recognize that LLMs’ effectiveness is primarily illustrated in language-related tasks. LLMs are not substitutes for the critical examination of findings or the nuanced and domain-specific understanding essential for truly bridging disciplinary language [21, 22].
Scientific collaboration has long been a compelling research topic within the academic community [23]. Previous studies have primarily used network analysis to characterize these collaborations [24, 25] or explored their effects on the quality and novelty of scientific outputs [26, 27]. Based on the evidence that LLMs are beginning to impact the way scholars work and collaborate, the questions then arise: Whether and how have LLMs transformed interdisciplinary collaboration? In what ways have they changed how different disciplines interact and collaborate? While a few recent studies have leveraged bibliometric analysis to examine scientific collaboration in the context of LLMs [28] or their applications in specific fields such as biomedical and health sciences [29], they do not address how LLMs have transformed interdisciplinary collaboration. In particular, given the general increase in team sizes and the rise of interdisciplinary research, as highlighted in prior studies [30, 31], demonstrating the unique impact of LLMs on scientific collaboration presents an additional challenge.
@@
To address the research gaps, our study explores the application of LLMs in scientific disciplines and their implications for interdisciplinary collaboration by addressing two research questions, as listed below. The first question examines the diversity of co-authors in terms of their institutional and departmental affiliations, using the Shannon Entropy measure. The second question aims to understand the structural patterns of co-authorship collaborations, using network analysis to identify key institutions and departments that are active producers of LLMs research, as well as those facilitating collaborations across disciplinary boundaries. To isolate the impact of LLMs from broader trends in scientific research, such as the general increase in team sizes or variations in productivity across different disciplines [31, 30, 32, 33], we compare the impact of LLMs against two control groups: papers from machine learning (ML papers) and papers that do not involve either LLM or ML (non-LLM/ML papers). This comparative approach allows us to assess whether the observed changes in collaboration patterns are attributable to the use of LLMs or if they are part of broader shifts occurring across all scientific disciplines.
@@
  • RQ1. How diverse are the coauthors of papers utilizing LLMs in terms of their research institutions and departments across different fields, in comparison to the control groups?
@@
  • RQ2. What are the structural patterns of co-authorship networks in research utilizing LLMs, and what roles do key entities (leading institutions and departments) play in facilitating and enhancing collaboration, in comparison to the control groups?
@@
To answer these questions, we collect 59,293 papers from OpenAlex related to LLMs from 27 scientific fields published between the release of the BERT model in October 2018 and September 2024. We use Shannon Entropy to measure the authorship diversity and network techniques to reveal the collaboration structures. Our findings indicate that since the advent of LLMs, particularly after the release of ChatGPT in November 2022, collaboration diversity has increased across many fields, notably Computer Science and Social Science. Overall, our findings demonstrate that LLM research has grown exponentially since 2022 and holds the potential to enhance interdisciplinary collaboration. It also encourages the involvement of prominent academic institutions in leading LLM-based research and applications in different domains.
@@ 2 Data and Methods
Aligned with the two research questions, our data collection and methods involve two units of analysis: papers and authors. The first research question examines papers to analyze the diversity of coauthors across research fields and institutions. The second research question centers on authors to explore the structural patterns of co-authorship networks and identify key researchers and their roles in facilitating collaboration. Combining these two units of analysis allows us to capture (1) which disciplines and institutions are most impacted by the advent of LLMs in terms of collaboration diversity, as well as (2) key researchers in LLM research and their respective departmental and institutional affiliations. These analyses help us understand how LLMs are impacting scientific collaboration by revealing patterns of interdisciplinary and cross-institutional partnerships, as well as identifying the influential actors driving LLM research across disciplinary boundaries. For each of the two sets of measures, we compare the results against two control groups (ML and non-LLM papers) to ensure the observed trends are specific to LLM researchand not reflective of broader patterns in other AI-related or unrelated fields.
@@ 2.1 Data preparation
The workflow for data cleaning is presented in subsection A.1. Within OpenAlex, we collect relevant papers on the most popular LLMs and their respective models, as detailed in Table 3 in subsection A.2. We use two general terms, “large language model” and “LLM,” along with popular open-source models (e.g., BERT, Flan-T5, LLaMA) and closed-source models (e.g., ChatGPT, Claude) based on the MMLU benchmark [35]. We observe that some models, like Yi or Phi, do not yield relevant papers, potentially introducing significant noise during paper screening. Additionally, models like grok-1 or Galactic did not return any search results. Our data collection results in a total of 177,462 papers.
To ensure the collected papers are relevant to the topic of LLMs, we restrict our search to titles and abstracts. However, some papers containing these keywords might still be irrelevant. Therefore, we implement several steps to filter out irrelevant papers. First, we consider only articles and preprints, excluding types such as editorials and opinions. Second, we remove potential duplicates, including those with duplicated titles and papers initially published as preprints and later as journal articles with similar titles. We use Jaccard similarity to identify and remove these duplicates (see subsection A.3). Third, we employ GPT-4o mini models to evaluate the relevance of a paper to LLM topics based on its title, abstract, and keywords, and filter out those papers that are not relevant to LLM topics (see subsection A.4). Applying these filtering criteria reduces the original collection to 59,293 papers for the subsequent analysis.
As mentioned earlier, an important question is whether the observed trend could be unique to LLM research or simply reflects a natural progression in academic collaboration. To investigate this, we establish two control groups. The first control group focuses on machine learning (ML) papers. We select ML as a control because it is a well-established field from which LLM emerged as a subfield. To construct this group, we collect a random sampling of 70,945 papers containing the phrase “machine learning” in either their title or abstract. Rather than searching for specific ML models like random forest or logistic regression, we use this broader criterion to gather a comprehensive sample while maintaining the simplicity and clarity of resulting dataset. To provide an even broader perspective beyond AI-related fields, we create a second control group consisting of a random sample of 73,110 papers from all other research categories—specifically, papers that belong neither to the ML nor LLM categories. This selection allows us to analyze collaboration patterns in the broader scientific community. The development of control groups is specifically documented in subsection A.1.
@@ 2.2 Measure of collaboration diversity
where $x_{i}$ denotes the $i$th author of a paper, $A(x_{i})$ denotes the set of authors’ affiliation information given a paper, $D(x_{i})$ represents their department information, $I(x_{i})$ represents their institution information, and $C(x_{i})$ represents their country information. It is important to note that these three types of information can vary significantly within a single paper. For instance, all collaborating authors could belong to the same institution and country but be affiliated with different departments. Our subsequent analysis particularly focuses on the collaboration between institutions and departments; therefore, we consider the first two sets of variables in Equation 1.
@@
While the entropy meassure provides valuable insights into collaboration diversity patterns over time, we are also interested in understanding how the launch of ChatGPT might have affected the scientific collaboration diversity. However, the entropy changes after the launch of ChatGPT cannot be simply attributed to causal effects due to the inherent trends in entropy over time. To further investigate the impact of ChatGPT on the collaboration diversity, a Difference-in-Difference (DiD) model is fitted for each field [36, 37] using the following equation:
$H(Y)=\beta_{0}+\beta_{1}\cdot\text{Treatment}+\beta_{2}\cdot\text{Post}+\beta_{3}\cdot(\text{Treatment}\times\text{Post})+e$ (3)
where $H(Y)$ is the entropy given the time; Treatment is a dummy variable indicating the treatment group (LLM); Post is a dummy variable indicating pre (=0) and post (=1) the launch of ChatGPT; Treatment $\times$ Post is a dummy variable indicating whether the outcome is observed in the treatment group and after the intervention. In addition, $\beta_{3}$ represents the change in the average entropy of the treatment group after the launch of ChatGPT, compared to what would have been expected in the absence of ChatGPT.
2.3 Measures of network structure and network comparison
We construct co-authorship networks for the LLM papers group, ML group, and non-LLM/ML group, based on the bipartite network projection, which involves converting a paper-author network to a co-authorship network whereby two authors are connected if they have co-authored at least one paper together. Each connection is weighted based on the total number of papers that each pair of researchers co-authored together. This method, as described by [38, 2], allows us to identify key researchers that have notable collaborative influence in the field, as well as any differences in collaboration patterns depending on the authors’ disciplines, institutions, and countries. The formula for our weighted projection approach is shown below:
@@ 2.3 Measures of network structure and network comparison
To measure structural similarity between the LLM group, ML group, and non-LLM/ML group networks, we conducted comparisons at the node, edge, and network levels. This approach has been validated in prior studies [40, 41], which found that comparing networks at all three levels because similarities can emerge from local-level interactions and extend to the overall network.
First, we measure node overlap using the normalized pairwise Jaccard index, which is a similarity measure based on the proportion of shared nodes between two networks relative to the total unique nodes in both networks. Second, edge overlap is also calculated using the Jaccard index applied to the adjacency matrices of the networks, so that we can determine the proportion of shared edges relative to the total unique edges. We require that the two networks share at least some common nodes [42], so edge overlap is only calculated when there is a non-zero node overlap.
Third, we compare network-level structural similarity using Netsimile [43], a graph embedding method that measures similarity using seven structural features derived from both node- and edge-level properties. Based on the algorithm, we generate a feature vector for each network based on these features: (1) degree, the number of edges connected to each node; (2) clustering coefficient, the proportion of a node’s neighbors that are connected to each other; (3) egonet size, the number of edges within a node’s neighborhood; (4) egonet density, the ratio of actual to possible edges in each node’s ego-network; (5) outward edges, the number of edges connecting the node to others outside its egonet; (6) average neighbor degree, the mean degree of a node’s neighbors; and (7) average neighbor clustering coefficient, the mean clustering coefficient of the node’s neighbors. We then compare the feature vectors between each pair of networks using Canberra distance, because it is effective at detecting small differences, especially for values near zero, and it normalizes differences to ensure that all features contribute fairly to the overall comparison [43]. The formula for this measure is below:
$NetSimileCanberra(G_{1},G_{2})=\sum_{i=1}^{n}\frac{\mid{G_{1_{i}}}-G_{2_{i}}\mid}{{|G_{1_{i}}|}+|G_{2_{i}}|}$ (5)
Here, $G_{1}$ and $G_{2}$ are feature vectors for the two networks, with $G_{1_{i}}$ and $G_{2_{i}}$ as their $i$th elements. The Canberra distance sums the normalized differences between elements, where smaller values indicate higher network similarity.
@@ 3 Results
In subsection 3.1, we analyze the diversity of collaboration based on the institutions and departments affiliated with authors in various fields. In subsection 3.2, we examine the author network to gain insight into how various institutions or departments collaborate, focusing particularly on comparing the collaboration patterns between LLM-related studies and those of control groups. Note that a paper’s “field“ categorization is provided by OpenAlex (see Appendix C for more details). OpenAlex uses an LLM, specifically a BERT model, to generate a score distribution for identified topics using a paper’s title, abstract, and citations [34]. The model provides up to three topics with the highest scores, which we map to fields according to the ASJC structure [34].
@@
We sort the collected 59,293 papers according to the count of papers in order, as presented in Figure 1(a). For the subsequent analysis, we focus on the top 12 fields with the most publications in the topics of LLMs: (1) Computer Science, (2) Medicine, (3) Social Science, (4) Engineering, (5) Decision Science, (6) Psychology, (7) Biochemistry, Genetics & Molecular Biology, (8) Business, Management & Accounting, (9) Health Professions, (10) Arts & Humanities, (11) Economics, Econometrics & Finance, and (12) Neuroscience. We then filter for papers with complete authors’ affiliation information, resulting in 30,075 papers with complete institution information and 22,453 papers with complete department information. Building on this approach, the ML group and the Non-LLM/ML group include 56,658 and 48,795 papers with institution information, respectively, and 47,002 and 40,579 papers with department information.
Several noteworthy observations emerge from the analysis. First, the Non-LLM/ML group demonstrates consistency in the number of papers, while the ML group exhibits a modest upward trajectory. In contrast, the LLM group shows a pronounced increase in the number of papers, especially after the release of ChatGPT (the dotted gray line in Figure 1(b)). Second, when examining the coefficient and r-value of the entropy, the LLM group reveals a steeper regression line. Specifically, the entropy increases based on authors’ affiliated institution ($coefficient=0.010$) and department ($coefficient=0.013$), which are significant compared to the two control groups.
@@
Figure 1: (a) The distribution of papers in the collection by field. (b) The temporal pattern of entropy and paper count based on authors’ affiliated institution (top) and department (bottom) information, respectively. In both subplots, the x-axis denotes the date by year. The primary y-axis denotes the Shannon Entropy calculated using Equation 2, while the secondary y-axis indicates the count of papers. Each point represents the averaged entropy based on papers published in one quarter. The dotted gray line indicates when OpenAI released ChatGPT.
3.1 Analysis and comparison of collaboration diversity
Figure 2(a) and Figure 2(b) show the collaboration diversity measured by Shannon Entropy for authors’ institutions and departments across three paper groups, respectively. Our analysis in this section addresses the following three key aspects.
1. What do the overall collaboration patterns in the LLM group reveal? Regarding the entropy analysis based on institutions and departments from the LLM group, the Shannon Entropy trends in Figure 2(a) and Figure 2(b) (blue line) closely align for each respective field. Most fields show a minor increase in the averaged entropy in 2024. In particular, fields like Computer Science, Medicine, Social Science, and Biochemistry, Genetics & Molecular Biology have shown consistent increases in institutional collaboration. Comparing different fields also reveals distinct patterns. Health-related fields, such as Medicine, Biochemistry, Genetics & Molecular Biology, Health Professions, and Neuroscience, demonstrate higher values in both institution and department entropy levels. This indicates more diverse collaboration patterns across these domains. Computer Science exhibits lower overall entropy, particularly in the early stage of LLM studies (e.g., BERT models), suggesting that LLM research initially remained concentrated within the computer science community. Additionally, Arts & Humanities displays notably lower institutional entropy compared to other disciplines.
However, the patterns of collaboration vary across disciplines and even between institutional and departmental entropy within the same field. For instance, in Engineering, institutional entropy remains relatively stable while departmental entropy shows high levels in early periods, suggesting that BERT-related research involved cross-departmental collaboration within the same institution. Medicine demonstrates a distinct pattern where institutional entropy shows an upward trend, while departmental entropy remains relatively constant. In contrast, Psychology and Biochemistry, Genetics & Molecular Biology exhibit reversed patterns compared to Medicine.
2. Do these collaboration patterns exhibit a general trend comparable to the control paper groups, or are they unique to the LLM paper group? We compare the entropy for each field across the three groups. Our analysis reveals several consistent patterns. The ML group consistently displays higher entropy levels than the other two paper groups; this is particularly pronounced in fields such as Medicine, Computer Science, Psychology, Social Science, and Biochemistry, Genetics & Molecular Biology. This likely reflects the well-established nature of research collaborations across disciplines in ML studies. However, when examining the slopes of the smoothing lines, the LLM group exhibits a more pronounced increasing trend. This is especially evident in Computer Science, Medicine, and Social Science, where the trend accelerates notably following ChatGPT’s release. Additionally, institution entropy remains relatively stable for both the ML and Non-LLM/ML groups, as compared to the LLM paper group.
Despite common trends, collaboration patterns vary significantly across disciplines. In Computer Science and Social Science, the LLM paper group demonstrates a more pronounced increasing trend in collaboration, whereas this pattern is less evident in Engineering and Decision Science. Health-related fields also exhibit distinct patterns: Medicine shows a marked increase in institutional collaboration within the LLM group, while department entropy remains similar across all three paper groups. In contrast, Psychology and Biochemistry, Genetics & Molecular Biology display clearer upward trends in department entropy for the LLM group compared to their institution entropy patterns.